OpenStudy (anonymous):

What is 21/35 in simplest form

OpenStudy (anonymous):

21/35=(7*3)/(7*5)=3/5

sammixboo (sammixboo):

First find the highest common factor both 31 and 35 have Factors of 21 = 1, 3, 7, 21 Factors of 35 = 1, 5, 7, 35 What is the highest number do they both have in common?

sammixboo (sammixboo):

21 and 35*

OpenStudy (dmndlife24):

Ohh no not this again